Modify Methionine, below, to show its zwitterion formModify Methionine, below, to show its zwitterion
Answer #1

General guidance

Concepts and reason

The neutral structure of amino acid, methionine, is given and the zwitter form needs to be drawn. This can be done by proton transfer from carboxylic acid to the amino group.

Fundamentals

The carboxylic acids containing an amino group (\u2212NH2)\\left( { - {\\rm{N}}{{\\rm{H}}_2}} \\right) are known as amino acids. The skeleton of an amino acid is as follows:

Side Chain R COOHCarboxylic group NH2 Amino group

Where R is an alkyl, aryl or heterocyclic group.

Zwitter ion: A zwitter ion is formed by self-neutralization reaction within an amino acid. The positive charge is equal to the negative charge in a zwitter ion and thus, the overall molecule is neutral and has no net charge.

Step-by-step

Step 1 of 2

Write the given structure of methionine and encircle the amino and carboxylic groups as shown below:

Explanation

The carboxylic and the amino groups in order to identify the points of proton transfer taking place in next step.

Step 2 of 2

The zwitter ion is as follows:

The zwitter ion of methionine is as follows:


Explanation

In the above step, proton transfer takes place from carboxylic group(COOH)\\left( {{\\rm{COOH}}} \\right) to the amino group(NH2)\\left( {{\\rm{N}}{{\\rm{H}}_2}} \\right) to form the corresponding zwitter ion of methionine. The zwitter ion formed is a dipolar molecule having equal positive and negative charges, making it overall a neutral molecule.

Answer

The zwitter ion of methionine is as follows: